# Parallel computations for Metropolis Markov chains with Picard maps

### v1

@@ Abstract
We develop parallel algorithms for simulating zeroth-order (aka gradient-free) Metropolis Markov chains based on the Picard map. For Random Walk Metropolis Markov chains targeting log-concave distributions $\pi$ on $\mathbb{R}^{d}$, our algorithm generates samples close to $\pi$ in $\mathcal{O}(\sqrt{d})$ parallel iterations with $\mathcal{O}(\sqrt{d})$ processors, therefore speeding up the convergence of the corresponding sequential implementation by a factor $\sqrt{d}$. Furthermore, a modification of our algorithm generates samples from an approximate measure $\pi_{\epsilon}$ in $\mathcal{O}(1)$ parallel iterations and $\mathcal{O}(d)$ processors. We empirically assess the performance of the proposed algorithms in high-dimensional regression problems and an epidemic model where the gradient is unavailable. Our algorithms are straightforward to implement and may constitute a useful tool for practitioners seeking to sample from a prescribed distribution $\pi$ using only point-wise evaluations of $\log\pi$ and parallel computing.
Key words— Parallel computing, Metropolis-Hastings, sampling, log-concavity, Gradient-free, Zeroth-order
@@ 1 Introduction
1.1 Picard map for parallel-in-time Markov chain simulation
@@
The Picard recursion is guaranteed to converge to its fixed point in at most $K$ iterations because, by construction, $X^{(j)}$ coincides with the solution of (1) in the first $j$ coordinates. However, since the sequential recursion in (1) also requires $K$ iterations to compute $X$, relevant applications of the Picard recursion are those where $X^{(j)}$ gets close (or exactly equal) to its fixed point for $j\ll K$ iterations, trading the use of parallel computing for a shorter run-time.
@@
Figure 1: Traces on the $(x_{1},x_{2})$-plane of the Picard recursion $X^{(j)}$ for $K=1000$ and $j=1,2,10,13$ (top left - bottom right). The time-index of each sequence is shown with a yellow-red gradient color. The underlying Markov chain is a $d=100$ dimensional Random Walk Metropolis with step-size $\xi=2/\sqrt{d}$ targeting a standard Gaussian distribution. The gray line is the fixed point (i.e. the output of the sequential algorithm). The dashed line corresponds to the part of the trajectory that has converged to its fixed point.
1.2 Zeroth-order Markov chain Monte Carlo
Markov chain Monte Carlo (MCMC) methods are the computational workhorse of many scientific areas, including Bayesian Statistics. These methods consist of simulating a Markov chain $(X_{0},X_{1},\dots)$ whose limiting measure coincides with a probability distribution of interest $\pi$, such as a Bayesian posterior. Then, ergodic averages $\frac{1}{N}\sum_{i=1}^{N}g(X_{i})$ can be used to estimate expectations $\int g(x)\pi(\mathrm{d}x)$ for any integrable function $g$, see e.g. Roberts and Rosenthal (2004). We are interested in the simulation of zeroth-order (aka gradient-free, derivative-free) Markov chains, which only require point-wise evaluations of $\log\pi$ (up to an additive unknown constant). These methods are commonly used when gradient information is unavailable, for either practical, computational, or mathematical reasons (cf. Conn et al., 2009, Chapter 1). This may occur for example in Bayesian applications with black-box likelihood evaluations (e.g. in form of proprietary code, code written by domain experts, or models requiring complex numerical solvers) or when the gradient is not well-defined such as in models with censored data (e.g. the SIR model considered in Section 6), in Pseudo-Marginal MCMC (Andrieu and Roberts, 2009) and in Approximate Bayes Computation (ABC) MCMC setting (Marjoram et al., 2003). See e.g. Conn et al. (2009) and other references on zeroth-order optimization for more discussion about applications of gradient-free methods.
Our objective is to leverage the use of $K>1$ parallel processors for reducing the parallel round complexity of this class of algorithms, defined as the number of point-wise evaluations of $\log\pi$ per parallel processor in order to obtain samples from $\pi$. Under standard log-concavity assumptions on $\pi$, the best known complexity result w.r.t. the number of dimensions $d$ for gradient-free sequential algorithms (i.e. $K=1$) is $\mathcal{O}(d)$ evaluations per sample, and it is achieved by classical Random Walk Metropolis [^1] Throughout the paper, the $\mathcal{O}$ notation ignores constants and logarithmic terms (Dwivedi et al., 2019; Andrieu et al., 2024). Note that this complexity matches the one of gradient-free convex optimization (Nesterov and Spokoiny, 2017).
1.3 Main results
Corollary 2 shows that, given a log-concave target density $\pi$ on $\mathbb{R}^{d}$ and $K=\mathcal{O}(\sqrt{d})$ processors, our proposed Online Picard algorithm applied to a Random Walk Metropolis Markov chain (initialized from a so-called feasible start) generates samples from a measure close to $\pi$ in total variation distance after $\mathcal{O}(\sqrt{d})$ parallel iterations. Thus, the Online Picard algorithm accelerates zeroth-order sequential algorithms by a factor of $\sqrt{d}$ with $K=\mathcal{O}(\sqrt{d})$ parallel processors, achieving optimal speedup (i.e. linear in $K$ up to logarithmic factors) for $K$ up to $\mathcal{O}(\sqrt{d})$. On the contrary, for larger values of $K$, we observe, both theoretically and empirically, a sharp decrease in relative gain. Motivated by this bottleneck and leveraging insights from the analysis of the Online Picard algorithm, we also propose an Approximate Picard algorithm which generates samples from an approximate measure $\pi_{\epsilon}$ after $\mathcal{O}(1)$ parallel iterations with $K=\mathcal{O}(d)$ processors. The error of the invariant measure $\pi_{\epsilon}$ relative to $\pi$ is assessed through numerical simulations, while a rigorous analysis lies beyond the scope of this work. Section 4 provides convergence results also for Picard algorithms applied to Metropolis within Gibbs Markov chains, which we empirically observe to perform better. The empirical performance observed in Section 6 closely matches the theory developed and shows substantial speedups (up to over 100 times) relative to the corresponding sequential algorithms, with highest efficiency obtained by Metropolis within Gibbs schemes. All proofs are provided in the appendix.
These results, combined with the simplicity of our algorithms, offer practitioners promising directions to parallelize computations for Bayesian problems with black-box, expensive models and no access to gradient information.
1.4 Related literature on parallel sampling and Picard maps
Previous attempts to parallelize computations for (zeroth-order) MCMC include running multiple independent chains (Rosenthal, 2000), which is straightforward to implement but does not reduce the burn-in or convergence period, or using pre-fetching (Brockwell, 2006) and multiple-try (Frenkel and Smit, 1996; Liu et al., 2000) algorithms, which however achieve complexity gains of only $\mathcal{O}(\log(K))$ in typical zeroth-order settings (Pozza and Zanella, 2024). The past year has seen a growing interest in parallel sampling based on Picard recursions applied to first-order (i.e. gradient-based) Markov chains for both Monte Carlo methods (Yu and Dalalyana, 2024; Anari et al., 2024) and generative models (Shih et al., 2023; Zhou and Sugiyama, 2024; Chen et al., 2024; Gupta et al., 2024). However, existing analyses focus on the (so-called unadjusted) setting where the function $f$ in (1) is smooth with respect to their first argument, which is not the case for Metropolis-Hastings algorithms – see (8) below. Our work avoids the classical smoothness conditions on $f$ and instead focuses on Metropolis-type Markov chains, where $f$ is piecewise constant with respect to its first argument. This setting leads to a substantially different analysis as well as convergence results for the induced Picard map, as highlighted in Section 2 and Section 3.4 below.
2 Picard algorithms with piecewise constant map $\Phi$
We consider Markov chains that satisfy the following assumption.
Assumption 1.
For every $W\in\mathcal{W}^{K}$, the Picard map $X\mapsto\Phi(X,W)$ is constant in a neighborhood of its fixed point.
@@ 2 Picard algorithms with piecewise constant map $\Phi$
Assumption 1 is satisfied by standard zeroth order Metropolis-Hastings kernels described in Section 3.2, which are our main motivation. Other relevant examples satisfying this assumption include Markov chains defined on discrete spaces. Assumption 1 has many implications. For example, it implies that the fixed point of $\Phi$ can be reached exactly after $j<K$ parallel iterations. When the fixed point is reached, the output of the Picard recursion matches that of the sequential counterpart, with no additional bias induced by the Picard recursion. This is in contrast to classical applications of Picard recursions, where stopping the Picard recursion before $K$ parallel iterations introduces bias and requires tuning of an additional error parameter.
2.1 Online Picard algorithm
@@
Assume we seek to simulate $N$ steps of the Markov chain defined in (1), $(X_{0},\dots,X_{N})$, using $K\leq N$ independent processors. In the classical Picard algorithm, one updates the first $K$ steps, $(X_{1},\dots,X_{K})$, with the Picard recursion until a stopping criterion is met, then proceeds with the next $K$ ones, $(X_{K+1},\dots,X_{2K})$, and so on until reaching $X_{N}$. Under Assumption 1, instead, one can monitor the coordinates that have already reached their fixed point while the algorithm runs, and avoid wasting computational power in updating them further, directly allocating available processors to subsequent coordinates. This leads to the following Online Picard algorithm: given a sequence of noise $W=(W_{0},W_{1},\dots)\in\mathcal{W}^{\infty}$ and initialization $X^{(0)}\in\mathcal{X}^{\infty}$, define the $j$-th parallel recursion of the Online Picard algorithm mapping $X^{(j)}\in\mathcal{X}^{\infty}$ to $X^{(j+1)}\in\mathcal{X}^{\infty}$ as
$X_{i}^{(j+1)}$ =$=\begin{cases}X_{i}^{(j)}&i\leq L^{(j)},\\ \Phi_{i-L^{(j)}}\left(X_{L^{(j)}:U^{(j)}}^{(j)},W_{L^{(j)}:U^{(j)}-1}\right)&L^{(j)}<i\leq U^{(j)}\,,\\ X_{U^{(j)}}^{(j+1)}&i>U^{(j)},\end{cases}$ (3)
@@
The Online Picard algorithm applies the Picard map to the indices $i\in\{L^{(j)},L^{(j)}+1,\dots,U^{(j)}\}$, where $L^{(j)}$ is an increasing sequence which monitors the position of the last index $i$ such that $X^{(j)}_{0:i}$ has reached its fixed point. The algorithm is then stopped as soon as $L^{(j)}\geq N$. While (3) is formally defined as a map on $\mathcal{X}^{\infty}$, the algorithm in practice only operates on $K$ coordinates of the Markov chain, as shown by Proposition 1 below and by the pseudocode provided in Appendix C. As for the Picard recursion in (2), the online recursion in (3) requires $U^{(j)}-L^{(j)}=K$ calls of the function $f$ that can be computed in parallel.
The Online Picard algorithm is always more efficient than the classical Picard algorithm which applies sequentially the Picard recursion of (2) until convergence of each subsequent block $\{X_{i}\colon\ell K<i\leq(\ell+1)K\}$ for $\ell=0,1,\dots$, since it avoids the vacuous updates of $X^{(j)}_{i}$ for $i\leq L^{(j)}$. See Figure 2 for an illustration.
Figure 2: Illustration of the Picard recursion applied sequentially to every block of length $K$ (left grid) vs the Online Picard algorithm (right grid). The color of the $(j,i)$ entry of each grid represents the state of the $i$th step of the Markov chain at the $j$th Picard recursion: red for $f(X^{(j)}_{i},W_{i})=f(X^{(j-1)}_{i},W_{i})$ (correct guess), blue for $f(X^{(j)}_{i},W_{i})\neq f(X^{(j-1)}_{i},W_{i})$. Black for the increments for which no processor has been allocated for computing the function $f$. Here, $K=4$, $N=3K$. Yellow boundary line in correspondence of $L^{(j)}=\sup\{i\leq U^{(j)}\colon X_{\ell}^{(j)}=X_{\ell}^{(j-1)};\ell\leq i\}$.
2.2 Markov structure of Online Picard algorithms
Let $(X^{(j)},W^{(j)},L^{(j)})_{j=0,1,\dots}$ be the sequence of random variables produced by the Online Picard recursion of (3)-(4). Define $\bar{X}^{(j)}=X^{(j)}_{L^{(j)}:U^{(j)}}$ and $\bar{W}^{(j)}=W_{L^{(j)}:U^{(j)}-1}$.
Proposition 1.
Proof.
$\bar{X}^{(j+1)}_{i}$ =$=\begin{cases}\Phi_{G^{(j)}+i}(\bar{X}^{(j)},\bar{W}^{(j)})&i\leq K-G^{(j)}\,,\\ \Phi_{K}(\bar{X}^{(j)}\,,\bar{W}^{(j)})&i>K-G^{(j)}\,,\\ \end{cases}$ $i=0,1,\dots,K\,,$ (5)
$\bar{W}^{(j+1)}_{i}$ =$=\begin{cases}\bar{W}^{(j)}_{G^{(j)}+i}&i\leq K-G^{(j)}-1\,,\\ \tilde{W}_{i,j}&i>K-G^{(j)}-1\,,\end{cases}$ $i=0,1,\dots,K-1\,,$ (6)
and (5)-(6) define the one-step Markov transition kernel of $(\bar{X}^{(j)},\bar{W}^{(j)})_{j=0,1,\dots}$. ∎
Proposition 1 sheds lights to the underlying (Markov) structure of Online Picard recursion. It can also be used for a memory-wise and computationally efficient implementation of the algorithm that iterates over a vector of length $K$. In the next section, we study the distribution of the random variables $(L^{(j)})_{j\in\mathbb{N}}$ for popular zeroth-order Markov chains under log-concavity, in order to derive parallel round complexity results for the Online Picard algorithm.
3 Complexity of Picard algorithms for zeroth-order Monte Carlo

### v2

@@ Abstract
We develop parallel algorithms for simulating zeroth-order (aka gradient-free) Metropolis Markov chains based on the Picard map. For Random Walk Metropolis Markov chains targeting log-concave distributions $\pi$ on $\mathbb{R}^{d}$, our algorithm generates samples close to $\pi$ in $\mathcal{O}(\sqrt{d})$ parallel iterations with $\mathcal{O}(\sqrt{d})$ processors, therefore speeding up the convergence of the corresponding sequential implementation by a factor $\sqrt{d}$. Furthermore, a modification of our algorithm generates samples from an approximate measure $\pi_{r}$ in $\mathcal{O}(1)$ parallel iterations and $\mathcal{O}(d)$ processors. We empirically assess the performance of the proposed algorithms in high-dimensional regression problems, an epidemic model where the gradient is unavailable and a real-word application in precision medicine. Our algorithms are straightforward to implement and may constitute a useful tool for practitioners seeking to sample from a prescribed distribution $\pi$ using only point-wise evaluations of $\log\pi$ and parallel computing.
keywords:
parallel computing, Metropolis-Hastings, sampling, log-concavity, gradient-free, zeroth-order
journal: Biometrika
@@ 1 Introduction
1.1 Zeroth-order Markov chain Monte Carlo
Markov chain Monte Carlo (MCMC) methods are the computational workhorse of many scientific areas. These methods consist of simulating a Markov chain $(X_{0},X_{1},\dots)$ whose limiting measure coincides with a probability distribution of interest $\pi$, such as a posterior distribution in Bayesian statistics. Then, ergodic averages $\frac{1}{N}\sum_{i=1}^{N}g(X_{i})$ can be used to estimate expectations $\int g(x)\pi(\mathrm{d}x)$ for any integrable function $g$, see e.g. Roberts & Rosenthal (2004).
In this paper we focus on zeroth-order (aka gradient-free or derivative-free) MCMC methods, which only require point-wise evaluations of $\log\pi$ up to an additive unknown constant, as opposed to first-order methods, which require computing the gradient of $\log\pi$. Zeroth-order methods are commonly used when gradient information is unavailable, for either practical, computational, or mathematical reasons (cf. Conn et al., 2009, Chapter 1). This may occur for example in Bayesian applications with black-box likelihood evaluations (e.g. in form of proprietary code, code written by domain experts, or models requiring complex numerical solvers) or when the gradient is not well-defined such as in models with censored data (e.g. the SIR model considered in Section 6), in Pseudo-Marginal MCMC (Andrieu & Roberts, 2009) and in Approximate Bayesian Computation (ABC) MCMC setting (Marjoram et al., 2003). See e.g. Conn et al. (2009); Carrillo et al. (2022); Pavliotis et al. (2022); Bou-Rabee et al. (2025); Grumitt et al. (2024) and other references on zeroth-order sampling and optimization for more discussion about applications of gradient-free methods.
For gradient-free sampling algorithms targeting log-concave distributions, the best complexity result w.r.t. the number of dimensions $d$ of the target $\pi$ is $\mathcal{O}(d)$, and it is achieved by classical Random Walk Metropolis (Dwivedi et al., 2019; Andrieu et al., 2024). Here and in what follows, the $\mathcal{O}$ notation ignores constants and logarithmic terms with respect to $d$ and complexity refers to the number of point-wise evaluations of $\log\pi$ required by the algorithm in order to obtain one sample from $\pi$ up to a fixed error. This complexity matches the one of gradient-free convex optimization (Nesterov & Spokoiny, 2017).
1.2 Parallel sampling
In this paper, we are interested in parallelization strategies for MCMC algorithms. This is particularly interesting given the increasing availability of parallel computing architectures – such as clusters of CPUs and Graphics Processing Units (GPUs) – which play an increasing central role in computation-based sciences. In the MCMC context, the simplest approach to parallelize computations is to run multiple independent chains. This is straightforward to implement in parallel but does not reduce the convergence period (also called burn-in or warm-up) of each chain (Rosenthal, 2000). Alternatively, one can consider algorithms which utilize $K>1$ parallel processors where, at each iteration, each processor evaluates independently and simultaneously $\log\pi$ for different parameter values and use the resulting information to speedup convergence to stationarity. Examples include pre-fetching methods (Brockwell, 2006), which compute $\log\pi$ at each potential state of the Markov chain for $j\geq 1$ steps ahead, and Multiple-try (Frenkel & Smit, 1996; Liu et al., 2000) which simulates $K$ proposal states at each iteration and computes $\log\pi$ in each proposed state before deciding which state to move to. Unfortunately, at least for zeroth-order methods and log-concave targets, these methods achieve only a $\mathcal{O}(\log(K))$ speedup factor, as recently proven in (Pozza & Zanella, 2024).
An alternative strategy to parallelize the simulation of Markov chains is provided by the Picard recursion detailed in Section 2.1 below, where the simulation of a Markov chain is reformulated as a fixed-point problem over trajectories. Methods based on the Picard recursion have recently witnessed a growing interest, particularly for first-order (i.e. gradient-based) Markov chains for both MCMC methods (Yu & Dalalyana, 2024; Anari et al., 2024) and generative models (Shih et al., 2023; Zhou & Sugiyama, 2024; Chen et al., 2024; Gupta et al., 2024). However, existing methods and analyses focus on first-order unadjusted settings where the function $f$ defined in (1) below is smooth with respect to its first argument, which is not the case for Metropolis-Hastings algorithms – see (5). Our work avoids the classical smoothness conditions on $f$ and instead focuses on zeroth-order Metropolis-type Markov chains, where $f$ is piecewise constant with respect to its first argument. This setting leads to a substantially different analysis and convergence results for the induced Picard map, as highlighted in Section 3.4 below.
1.3 Structure of the paper and main results
Section 2 introduces the Picard map and our proposed Online Picard algorithm. Section 3 presents its theoretical analysis. In particular, Corollary 3.6 shows that, given a log-concave target density $\pi$ on $\mathbb{R}^{d}$ and $K$ processors, with $K$ up to $\mathcal{O}(\sqrt{d})$, the Online Picard algorithm applied to a Random Walk Metropolis Markov chain obtains a sample from $\pi$ in $\mathcal{O}(d/K)$ iterations, thus accelerating zeroth-order sequential algorithms by a factor of $K$ and achieving optimal speedup (i.e. linear in $K$). To the best of our knowledge, this is the first parallel zeroth-order MCMC scheme with provably linear speedup in the canonical log-concave set-up. Section 4 extends the convergence results to Picard algorithms applied to Metropolis within Gibbs Markov chains, which we empirically observe to perform better. Section 5 proposes an Approximate Online Picard algorithm that scales well also when $K\gg\sqrt{d}$, at the price of introducing a bias in the invariant distribution, whose size is assessed through numerical simulations. For both the approximate and exact schemes, the empirical performance observed in Section 6 closely matches the theory developed and shows substantial speedups (up to over 100 times) relative to the corresponding sequential algorithms. All proofs are provided in the appendix.
These results, combined with the simplicity of the proposed algorithms, offer practitioners promising directions to parallelize computations for Bayesian problems with black-box, expensive models and no access to gradient information.
2 Sampling algorithms based on Picard maps
2.1 Picard map for Markov chain simulation
@@
The Picard recursion is guaranteed to converge to its fixed point in at most $K$ iterations because, by construction, $X^{(j)}$ coincides with the solution of (1) in the first $j$ coordinates. However, since the sequential recursion in (1) also requires $K$ iterations to compute $X$, relevant applications of the Picard recursion are those where $X^{(j)}$ gets close (or exactly equal) to its fixed point for $j\ll K$ iterations, thus exploiting parallel computing to shorten the overall runtime.
@@
Figure 1: Traces on the $(x_{1},x_{2})$-plane of the Picard recursion $X^{(j)}$ for $K=1000$ and $j=1,2,10,13$ (top left - bottom right). The time-index of each sequence is shown with a yellow-red gradient color. The underlying Markov chain is a $d=100$ dimensional Random Walk Metropolis with stepsize $\xi=2/\sqrt{d}$ targeting a standard Gaussian distribution. The gray line is the fixed point (i.e. the output of the sequential algorithm). The dashed line corresponds to the part of the trajectory that has converged to its fixed point.
2.2 Picard algorithms with piecewise constant map $\Phi$
@@
The Picard map applied to standard zeroth-order Markov chains (such as those described in Section 3.2) differs from those of classical applications because, for every noise variable $W\in\mathcal{W}^{K}$, the Picard map $X\mapsto\Phi(X,W)$ is piecewise constant. Other relevant examples satisfying this condition include Markov chains defined on discrete spaces. A piecewise constant map $\Phi$ has many implications. For example, it implies that the fixed point of $\Phi$ can be reached exactly after $j<K$ parallel iterations. When the fixed point is reached, the output of the Picard recursion matches that of the sequential counterpart, with no additional bias induced by the Picard recursion. This is in contrast to classical applications of Picard recursions, where stopping the Picard recursion before $K$ parallel iterations introduces bias and requires tuning of an additional error parameter. In what follows, we exploit this property to develop parallel algorithms that allocate computational resources more efficiently.
As above, assume we seek to simulate $N$ steps of the Markov chain defined in (1), $(X_{0},\dots,X_{N})$, using $K\leq N$ independent processors. In the classical Picard algorithm, one updates the first $K$ steps, $(X_{1},\dots,X_{K})$, with the Picard recursion in (2) until a stopping criterion is met, then proceeds with the next $K$ ones, $(X_{K+1},\dots,X_{2K})$, and so on until reaching $X_{N}$. See Algorithm 1 for the associated pseudocode in the case of piecewise constant maps.
@@
Alternatively, one can monitor the coordinates that have already reached their fixed point while the algorithm runs, and avoid wasting computational power in updating them further, directly allocating available processors to subsequent coordinates. This leads to the following Online Picard algorithm: given a sequence of noise $W=(W_{0},W_{1},\dots)\in\mathcal{W}^{\infty}$ and initialization $X^{(0)}\in\mathcal{X}^{\infty}$, define the $j$-th parallel recursion of the Online Picard algorithm mapping $X^{(j)}\in\mathcal{X}^{\infty}$ to $X^{(j+1)}\in\mathcal{X}^{\infty}$ as
$X_{i}^{(j+1)}$ $=\begin{cases}X_{i}^{(j)}&i\leq L^{(j)},\\ \Phi_{i-L^{(j)}}\left(X_{L^{(j)}:U^{(j)}}^{(j)},W_{L^{(j)}:U^{(j)}-1}\right)&L^{(j)}<i\leq U^{(j)}\,,\\ X_{U^{(j)}}^{(j+1)}&i>U^{(j)},\end{cases}$ (3)
@@
The Online Picard algorithm applies the Picard map to the indices $i\in\{L^{(j)},L^{(j)}+1,\dots,U^{(j)}\}$, where $L^{(j)}$ is an increasing sequence which monitors the position of the last index $i$ such that $X^{(j)}_{0:i}$ has reached its fixed point. The algorithm is then stopped as soon as $L^{(j)}\geq N$, see also the pseudocode in Algorithm 2. While (3) is formally defined as a map on $\mathcal{X}^{\infty}$, the algorithm in practice only operates on $K$ coordinates of the Markov chain as detailed in Appendix D in the supplementary material. The Online Picard algorithm is always more efficient than the classical Picard algorithm (Algorithm 1), since it avoids the vacuous updates of $X^{(j)}_{i}$ for $i\leq L^{(j)}$, see Figure 2 for an illustration.
Input: $N,K\in\mathbb{N}$, $X^{(0)}_{0:K}\in\mathcal{X}^{K+1}$.
Initialize $j=0$, $\ell=0$ and $W_{0},W_{1},\dots,W_{N-1}\overset{\text{i.i.d.}}{\sim}\nu$;
while $\ell K<N$ do
 $X^{(j+1)}_{\ell K:(\ell+1)K}=\Phi(X^{(j)}_{\ell K:(\ell+1)K},W_{\ell K:[(\ell+1)K-1]})$; This step uses parallel processors;
 $X_{i}^{(j+1)}=X^{(j)}_{i}$, for $i<\ell K$;
 if $X^{(j+1)}_{\ell K:(\ell+1)K}=X^{(j)}_{\ell K:(\ell+1)K}$ then
    $\ell=\ell+1$;
 $j=j+1$;
Output: $X^{(j)}_{0:N}$.
Algorithm 1 Picard algorithm.
Input: $N,K\in\mathbb{N}$, $X^{(0)}_{0:K}\in\mathcal{X}^{K+1}$
Initialize $j=0$, $L^{(0)}=0$ and $W_{0},W_{1},\dots,W_{N-1}\overset{\text{i.i.d.}}{\sim}\nu$;
while $L^{(j)}<N$ do
 $U^{(j)}=L^{(j)}+K$;
 $X^{(j+1)}_{L^{(j)}:U^{(j)}}=\Phi(X^{(j)}_{L^{(j)}:U^{(j)}},W_{L^{(j)}:U^{(j)}-1})$; This step uses parallel processors;
 $L^{(j+1)}=\sup\{i\leq U^{(j)}\colon f(X^{(j+1)}_{\ell},W_{\ell})=f(X^{(j)}_{\ell},W_{\ell})\text{ for }0\leq\ell<i\}$;
 $X_{i}^{(j+1)}=X^{(j+1)}_{U^{(j)}}$, for $i>U^{(j)}$, $X_{i}^{(j+1)}=X^{(j)}_{i}$, for $i\leq L^{(j)}$;
 $j=j+1$;
Output: $X^{(j)}_{0:N}$.
Algorithm 2 Online Picard algorithm
Figure 2: Illustration of the classical Picard algorithm (Algorithm 1) applied sequentially to every block of length $K$ (left grid) vs. the Online Picard algorithm (Algorithm 2, right grid). The color of the $(j,i)$ entry of each grid represents the state of the $i$th step of the Markov chain at the $j$th Picard recursion: red for $f(X^{(j)}_{i},W_{i})=f(X^{(j-1)}_{i},W_{i})$ (correct guess), blue for $f(X^{(j)}_{i},W_{i})\neq f(X^{(j-1)}_{i},W_{i})$. Black for the increments for which no processor has been allocated for computing the function $f$. Here, $K=4$, $N=3K$. Yellow boundary line in correspondence of $L^{(j)}=\sup\{i\leq U^{(j)}\colon X_{\ell}^{(j)}=X_{\ell}^{(j-1)};\ell\leq i\}$.
In the next two sections, we study the distribution of the random variables $(L^{(j)})_{j\in\mathbb{N}}$ for the Online Picard algorithm applied to popular zeroth-order Markov chains under log-concavity, which allows to deduce complexity statements about the algorithm in that setting.
3 Complexity results under log-concavity